# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ 3.4 Dipole transitions
$\left\langle\hat{q}_{i}\hat{q}_{j}\right\rangle_{\hat{q}}=\frac{1}{3}\delta_{ij}$ $\quad\text{for scalar emission},$ (3.17)
$\left\langle\sum_{\lambda}e^{(\lambda)}_{i}e^{(\lambda)}_{j}\right\rangle_{\hat{q}}=\frac{2}{3}\delta_{ij}$ $\quad\text{for vector emission}.$ (3.18)
Here, $\hat{q}_{i}$ and $e_{i}$ are the Cartesian components of the emission direction $\hat{q}=\mathbf{q}/|\mathbf{q}|$ and of the transverse polarization vectors $\mathbf{e}^{(\lambda)}$ in Coulomb gauge; the polarization sum only needs to run over the two transverse polarizations, as the emission of the longitudinal mode of $V$ is suppressed as $m_{V}^{2}/\omega^{2}$ (see below.)
@@
Below, the factor in parentheses is denoted by $C_{\phi}$. This result is in agreement in the case of emission from Dirac fermions $\chi$ with [71, 72], where the dipole amplitude was derived for electron-nucleon scalar bremsstrahlung in the $m_{e}\gg m_{p}$ limit; $m_{e(p)}$ is the electron (proton) mass. For massive gauge vector emission, taking into account the longitudinal part of the polarization vector $V_{\mu}$, the NR leading order emission amplitude becomes
@@
We point out that the $m_{\phi,V}$-dependent terms entering (3.20) and (3.21), prior to squaring, averaging over the emitted particle direction, and—in the latter—summing over polarizations, possess the same analytic structure for $\phi$-emission and for the longitudinal mode of $V$-emission. This results from the combined effects of Ward identity and NR reduction of emission amplitudes; see App. F for more details. The differences in the energy-differential cross-sections between scalar and vector emission are then captured in overall factors $C_{\phi}$ and $C_{V}$, respectively. Given a fraction $x\equiv 2\omega/\mu v_{i}^{2}\leq 1$ of CM energy carried away by the emitted particle, the cross-section for dipole emission then reads,
@@
Figure 2: Energy differential cross-section for dipole emission in distinguishable particle scattering. The solid brown lines correspond to the massless emission. The dotted colored lines show massive emission with $m_{\phi,V}/\omega_{\text{max}}=0.01$ (leftmost) and $m_{\phi,V}/\omega_{\text{max}}=0.1$ (rightmost). Black stars indicate values of the energy differential cross-section obtained from a fully numerical integration of unexpanded amplitudes. Left panel: Dipole, long-range mediation (3.22). Numerical integration parameters: $m_{2}=1$ MeV, $m_{1}=1$ GeV, and $v_{i}\approx 4\times 10^{-3}$. Right panel: Dipole, short-range mediation (3.26). Numerical integration parameters: $m_{2}=1$ MeV, $m_{1}=1$ GeV, $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV and $v_{i}\approx 4\times 10^{-3}$
@@
Similarly to the case of long-range mediation, we find back a model-independence of the kinematic structure in the emission amplitude. In the current scenario, where the mediator and the emitted particle differ, primed coupling constants $(\prime)$ are introduced that describe the mutual interaction of the colliding DM pair. Therefore, the modified dipole emission factors $\mathcal{F_{D}}$ for fermionic DM candidates now read
@@
Corrections to the squared amplitude for massive particle emission are identical to (3.20) and (3.21), and the factors $C_{\phi,V}$ remain unaltered. The energy differential dipole cross-section in the short-range mediated force regime is then given by
@@ 3.5.1 Scattering of distinguishable particles
We start by discussing quadruopole radiation for distinguishable DM particle scattering, i.e., the limit of canceling coupling-to-mass ratios of the DM pair.
@@
$\times\left\{\alpha\left[|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}+|\vec{v}_{i}|^{2}|\vec{v}_{f}|^{2}\left(2-4(\vec{v}_{i}\cdot\vec{v}_{f})^{2}\right)\right]+\beta\left(|\vec{v}_{i}|^{2}-|\vec{v}_{f}|^{2}\right)^{2}\right\}.$ (3.27)
Here, $(\alpha,\beta)$ are model-dependent numerical coefficients that are provided below. The squared amplitudes are obtained by taking the transverse polarization sum in the case of vector boson emission. The averaging over the direction of the emitted particle involves the identities
$\Big{\langle}\sum_{\lambda}{\hat{q}}_{j}{\hat{q}}_{k}{e}^{(\lambda)}_{l}{e}^{(\lambda)}_{m}\Big{\rangle}_{\bm{\hat{}}{\mathbf{q}}}$ $=\frac{1}{15}\left(-\delta_{jm}\delta_{kl}+4\delta_{jk}\delta_{lm}-\delta_{jl}\delta_{km}\right)\quad$ for vector emission, (3.28)
$\left\langle{\hat{q}}_{j}{\hat{q}}_{k}{\hat{q}}_{l}{\hat{q}}_{m}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}$ $=\frac{1}{15}\left(\delta_{jm}\delta_{kl}+\delta_{jk}\delta_{lm}+\delta_{jl}\delta_{km}\right)\quad$ for scalar emission.
At quadrupole order, the universality of emission amplitudes in the NR expansion across the various cases is now lost, so that the additional coefficients $\alpha$ and $\beta$ parameterizing the differences are introduced. This stems from differing NR currents for scalar and vector emission from fermions at this order in the $v_{i}$-expansion and a lost symmetry under the exchange of $\mathbf{q}$ and $\mathbf{e}$; see the respective equations D.6 and D.14 in App. D. The quadrupole emission factors, together with those coefficients, read
@@
It is important to emphasize that this contribution, although scaling as $\mathcal{O}(v_{i}^{-4})$, does not originate from quadrupole emission. Its overall coupling dependence is reduced by two powers compared to the latter. The associated energy-differential cross-section reads
@@
Expressions for the full amplitudes and differential cross sections that are not in any limit of $A_{j}/m_{j}$ relative to $\lambda_{j}$ are given in App. C.
As in the case of dipole emission, we may take into account the emitted particle mass $m_{\phi,V}$. In contrast to (3.20) and (3.21), there is no universal mass-dependent factor that could be pulled out in front of an amplitude of vanishing emitted particle mass. However, prior to squaring the full amplitude, the $m_{\phi,V}$-dependent terms in $\phi$-emission and in the longitudinal mode of $V$-emission agree; see App. F for a detailed discussion. We further note that in the case of massive scalar emission $\phi$, in the limit $(A_{j}/m_{j})^{2}\ll\lambda_{j}$, the NR expansion of the emission amplitude does not gain any dependency on $m_{\phi}$, so that any mass-dependent corrections are solely due to phase space in the cross-section.
In the left panel of Fig. 3, we plot the energy-differential cross sections given in (3.30) and (3.32). The mass-dependent contribution of the emitted particle, derived in App. F, is included and compared with the numerical integration of the full, unexpanded amplitude. To perform the comparison, we consider a system of distinguishable particles with equal coupling-to-mass ratios, ensuring that quadrupole emission provides the dominant contribution to the energy differential cross-section. Identical parameter values are used in both cases prior to any normalization, with fixed values of $v_{i},m_{j},m_{\phi,V}$, and the coupling constant $g_{j},y_{j},A_{j}/m_{j}$ and $\lambda_{j}$. The numerical results are in agreement with the analytical expressions, thereby validating the latter.
@@
Turning now to the case of a short-range interaction mediated by $\phi^{\prime}$ or $V^{\prime}$, the quadrupole emission amplitudes remain model-dependent. In the limits $(r\ll 1,\,m_{\phi^{\prime},V^{\prime}}<\sqrt{s})$, the squared, averaged over the initials spins, summed over the final spins and averaged over the direction of the emitted particles amplitudes are obtained from (3.5.1) with the replacement $|\vec{v}_{i}-\vec{v}_{f}|^{4}\rightarrow m_{\phi^{\prime},V^{\prime}}^{4}/\mu^{4}$ and using the following quadrupole emission factors and coefficients for fermionic DM candidates instead,
@@
These emission amplitudes lead to the energy differential cross-sections
@@
In the case of scalar emission from scalar DM candidates, when the emission is vertex-dominated (see App. C), the squared emission amplitude is obtained from (3.31) via the substitutions $|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}\rightarrow m_{\phi^{\prime},V^{\prime}}^{4}/\mu^{4}$, $\lambda_{j}\to\lambda^{\prime}_{j}$, and $A_{j}\to A^{\prime}_{j}$. Similarly to the long-range mediation case, this squared emission amplitude does not originate from quadrupole emission, despite exhibiting the same $v_{i}$-dependence. The associated energy differential cross-section is
@@
We note that, in the contact limit, this energy differential cross-section depends only on the primed coupling of the DM to the short-range mediator and on an interaction involving both mediators simultaneously. In contrast, the quadrupole energy differential cross-sections (3.36) receive contributions from the primed couplings of the DM to the short-range mediator as well as from the unprimed couplings to the long-range mediator. For massive emission, the agreement between the $m_{\phi,V}$-dependent term of the $\phi$-emission amplitude and the longitudinal component of $V$-emission prior squaring again holds; see App. F for a comprehensive treatment. For (3.37) the $m_{\phi}$-dependent correction appears from the phase space.
In the left panel of Fig. 4, we show the energy-differential cross sections given in (3.36) and (3.37), including the mass-dependent contribution of the emitted particle, as derived in App. F. The latter are compared to a numerical integration of the full, unexpanded amplitude. To enable a direct comparison, we consider a system of distinguishable particles with equal coupling-to-mass ratios, for which quadrupole emission constitutes the leading contribution to the energy-differential cross section. Identical values of $v_{i}$, $m_{j}$, $m_{\phi,V}$, $m_{\phi^{\prime},V^{\prime}}$, and the coupling constants [^8] We use the following shorthand notation in the case of emission process involving scalar DM candidate:   $C_{g^{\prime}_{j},\lambda_{S}}=g_{1}^{\prime}g^{\prime}_{2}+\lambda_{S}m^{2}_{V^{\prime}}/(4m_{1}m_{2})$, and $C_{A_{j}^{\prime},\lambda_{S}}=(A_{1}^{\prime}/m_{1})(A^{{}^{\prime}}_{2}/m_{2})-\lambda_{S}m^{2}_{\phi^{\prime}}/(2m_{S}^{2})$. $g^{(\prime)}_{j}$, $y^{(\prime)}_{j}$, $A^{(\prime)}_{j}/m_{j}$, and $\lambda_{S}$ are used in both cases, prior to any normalization. We find excellent agreement between the analytical and numerical results.
@@ 3.5.2 Scattering of identical particles
For the emission of a vector boson $V^{\mu}$, in contrast to the scattering of distinguishable particles, the emission amplitudes of the dSIDM scenarios $\chi\chi\to\chi\chi V$ and $SS\to SSV$ are not identical. This difference arises from the statistical properties of the DM candidates, which determine the interference sign between the contributions from the $t$- and $u$-channel diagrams. For vector emission from Dirac fermions, $\chi\chi\to\chi\chi V$, the energy differential cross-section reads
@@
We now turn to the emission of $\phi$ particles from fermions and scalars. For scalar emission from fermions, both Dirac fermions, $\chi\chi\to\chi\chi\phi$ and Majorana fermions, $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$, share the same energy differential cross-section
@@
Here, we are in disagreement with the results of [72] where scalar emission from electron-electron and nucleon-nucleon bremsstrahlung is derived. We ascertain our analytical result by comparing it against an exact numerical evaluation of the full 2-to-3 cross-section.
The energy differential cross-sections for $\phi$ emission from complex scalars, $SS\to SS\phi$, and real scalars $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$ are identical. In the limit $({A}/{m_{S}})^{2}\gg\lambda$, the external leg emission diagrams dominate, and the associated energy differential cross-section reads
@@
As a result, for fermionic DM candidates, the energy differential cross-section for massless emission is readily obtained by multiplying the right-hand side (RHS) of (3.36) by the factor $S_{f}=1/2$, accounting for the symmetry of identical particles in the final state, and by performing the following substitutions: $g^{(\prime)}_{j}\to g^{(\prime)}$, $y^{(\prime)}_{j}\to y^{(\prime)}$, and $m_{j}\to m_{\chi}$. In the case of massive emission, these changes must be applied in (F.19) for the process $\chi\chi\to\chi\chi V$, and in (F.22) for the processes $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$ and $\chi\chi\to\chi\chi\phi$.
For scalar DM candidates, the inclusion of the $u$-channel doubles the contribution of the external leg and vertex diagrams to the emission amplitude compared to the contact emission diagrams. As a result, the dissipative process $SS\to SSV$ has the following energy differential cross-section
@@
The scenarios involving the emission of a scalar boson from real or complex scalar DM candidates, i.e., $SS\to SS\phi$, and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$, share identical energy differential cross-sections across the two limiting cases previously defined in Sec. 3.2. For external leg-dominated emission (i.e., $\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}\gg\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$; see App. C for details), we find
@@
For vertex-dominated emission (see App. C) in the scenarios $SS\to SS\phi$, and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$, the energy differential cross-section is
@@
Figure 4: Quadrupole energy-differential cross section for the different dSIDM scenarios mediated by a short-range force. The massless emission limits are shown as solid colored lines, while the dotted colored line represents the massive emission with $m_{\phi,V}/\omega_{\text{max}}=0.1$. Orange stars indicate values of the energy differential cross-section obtained from a fully numerical integration of unexpanded amplitudes. The relevant coupling combinations $\{a,a^{\prime},a_{j}^{\prime}\}$ vary by scenario: i) for $\chi\chi\to\chi\chi V$, $\{g,g^{\prime},g_{j}^{\prime}\}$; ii) for $SS\to SSV$, $\{g,C^{1/2}_{g^{\prime},\lambda_{S}},C^{1/2}_{g_{j}^{\prime},\lambda_{S}}\}$; iii) for scalar emission from $\chi$($\tilde{\chi}$), $\{y,y^{\prime},y_{j}^{\prime}\}$; iv) for scalar emission with SS($\tilde{S}$) iv.a) in the external-leg regime (see App. C), $\{A/m_{S},C^{1/2}_{A^{\prime},\lambda_{S}},C^{1/2}_{A^{\prime}_{j},\lambda_{S}}\}$, iv.b) in the vertex-dominated regime (see App. C), $(a/m_{S,\chi})^{2}a^{\prime 2}_{1}a^{\prime 2}_{2}\to((A^{\prime}/m_{S})/m_{S})^{2}(\lambda^{\prime}_{1}+\lambda^{\prime}_{2})^{2}$ and $a^{\prime 4}a^{2}\to(A^{\prime}/m_{S})^{2}\lambda^{\prime 2}$. In the numerical analysis, these two regimes are represented by setting $A^{\prime}/m_{S}=100\,\lambda_{j}^{\prime}$ and $A^{(\prime)}/m_{S}=0.01\,\lambda_{j}^{\prime}$, respectively. Additionally, we assume $\lambda_{S}\sim g^{\prime}_{j},A^{\prime}_{j}/m_{j}$. Left panel: Distinguishable particle scattering (massless emission in Sec. 3.5.1 and massive emission in App. F) with equal coupling-to-mass ratio $a/m$. Numerical integration parameters: $m_{1}=0.1$ GeV, $m_{2}=1$ GeV, $a_{1}=0.1\,a_{2}$, $v_{i}\approx 4\times 10^{-3}$, and $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV. Right panel: Identical particle scattering (Sec. 3.5.2). Numerical integration parameters: $m_{S,\chi}=1$ GeV, $v_{i}\approx 6\times 10^{-3}$, and $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV.

### v2

@@ 3.4 Dipole transitions
$\left\langle\hat{q}^{i}\hat{q}^{j}\right\rangle_{\hat{q}}=\frac{1}{3}\delta^{ij}$ $\quad\text{for scalar emission},$ (3.23)
$\left\langle\sum_{\lambda}e_{(\lambda)}^{i}e_{(\lambda)}^{*j}\right\rangle_{\hat{q}}=\frac{2}{3}\delta^{ij}$ $\quad\text{for vector emission}.$ (3.24)
Here, $\hat{q}^{i}$ and $e^{*j}$ are the Cartesian components of the emission direction $\hat{q}=\mathbf{q}/|\mathbf{q}|$ and of the transverse polarization vectors $\mathbf{e}_{(\lambda)}^{*}$ in Coulomb gauge; the polarization sum only needs to run over the two transverse polarizations, as the emission of the longitudinal mode of $V$ is suppressed as $m_{V}^{2}/\omega^{2}$ (see below.)
@@
Below, the factor in parentheses is denoted by $C_{\phi}$. This result is in agreement in the case of emission from Dirac fermions $\chi$ with [82, 83], where the dipole amplitude was derived for electron-nucleon scalar bremsstrahlung in the $m_{e}\gg m_{p}$ limit; $m_{e(p)}$ is the electron (proton) mass. For massive gauge vector emission, taking into account the longitudinal part of the polarization vector $V^{\mu}$, the NR leading order emission amplitude becomes
@@
We point out that the $m_{\phi,V}$-dependent terms entering (3.26) and (3.27), prior to squaring, averaging over the emitted particle direction, and—in the latter—summing over polarizations, possess the same analytic structure for $\phi$-emission and for the longitudinal mode of $V$-emission. This results from the combined effects of Ward identity and NR reduction of emission amplitudes; see App. F for more details. The differences in the energy-differential cross sections between scalar and vector emission are then captured in overall factors $C_{\phi}$ and $C_{V}$, respectively. Given a fraction $x\equiv 2\omega/\mu v_{i}^{2}\leq 1$ of CM energy carried away by the emitted particle, the energy-differential cross section for dipole emission then reads,
@@
Figure 2: Energy-differential cross section for dipole emission in distinguishable particle scattering. The solid brown lines correspond to massless emission. The dotted colored lines show massive emission with $m_{\phi,V}/\omega_{\text{max}}=0.01$ (leftmost) and $m_{\phi,V}/\omega_{\text{max}}=0.1$ (rightmost). The black stars indicate values of the energy-differential cross section obtained from a fully numerical integration of unexpanded amplitudes. Left panel: Dipole, long-range mediation (3.28). Numerical integration parameters: $m_{2}=1$ MeV, $m_{1}=1$ GeV, and $v_{i}\approx 4\times 10^{-3}$. Right panel: Dipole, short-range mediation (3.32). Numerical integration parameters: $m_{2}=1$ MeV, $m_{1}=1$ GeV, $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV and $v_{i}\approx 4\times 10^{-3}$.
@@
Similarly to the case of long-range mediation, we find that the kinematic part of the dipole emission amplitude is model-independent. In the current scenario, where the mediator and the emitted particle differ, primed coupling constants $(\prime)$ are introduced that describe the mutual interaction of the colliding DM pair. Therefore, the modified dipole emission factors $\mathcal{F_{D}}$ for fermionic DM candidates now read
@@
Corrections to the squared amplitude for massive particle emission are identical to (3.26) and (3.27), and the factors $C_{\phi,V}$ remain unaltered. The dipole energy-differential cross section in the short-range mediated force regime is then given by
@@ 3.5.1 Scattering of distinguishable particles
We start by discussing quadrupole radiation for distinguishable DM particle scattering, i.e., the limit of canceling coupling-to-mass ratios of the DM pair.
@@
$\times\left\{\alpha\left[|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}+|\vec{v}_{i}|^{2}|\vec{v}_{f}|^{2}\left(2-4(\hat{\vec{v}}_{i}\cdot\hat{\vec{v}}_{f})^{2}\right)\right]+\beta\left(|\vec{v}_{i}|^{2}-|\vec{v}_{f}|^{2}\right)^{2}\right\}.$ (3.33)
Here, $\hat{v}_{i,f}=\mathbf{v}_{i,f}/|\mathbf{v}_{i,f}|$ and $(\alpha,\beta)$ are model-dependent numerical coefficients that are provided below. The squared amplitudes are obtained by taking the transverse polarization sum in the case of vector boson emission. The averaging over the direction of the emitted particle involves the identities given in (3.13) and (3.17).
At quadrupole order, the universality of emission amplitudes in the NR expansion across the various cases is now lost, so that the additional coefficients $\alpha$ and $\beta$ parameterizing the differences are introduced. The model dependence stems from differing NR currents for scalar and vector emission from fermions at this order in the $v_{i}$-expansion and a loss of symmetry under the exchange of $\mathbf{q}$ and $\mathbf{e}^{*}$; see the respective equations D.6 and D.14 in App. D. The quadrupole emission factors, together with those coefficients, read
@@
It is important to emphasize that this contribution, although scaling as $\mathcal{O}(v_{i}^{-4})$, does not originate from quadrupole emission. Its overall coupling dependence is reduced by two powers compared to the latter. The associated energy-differential cross section reads
@@
Expressions for the full amplitudes and differential cross sections that are not in any limit of $(A_{j}/m_{j})^{2}$ relative to $\lambda_{j}$ are given in App. C.
As in the case of dipole emission, we may take into account the emitted particle mass $m_{\phi,V}$. In contrast to (3.26) and (3.27), there is no universal mass-dependent factor that could be pulled out in front of an amplitude of vanishing emitted particle mass. However, prior to squaring the full amplitude, the $m_{\phi,V}$-dependent terms in $\phi$-emission and in the longitudinal mode of $V$-emission agree; see App. F for a detailed discussion. We further note that in the case of massive scalar emission $\phi$, in the limit $(A_{j}/m_{j})^{2}\ll\lambda_{j}$, the NR expansion of the emission amplitude does not gain any dependency on $m_{\phi}$, so that any mass-dependent corrections are solely due to phase space in the cross section.
In the left panel of Fig. 3, we plot the energy-differential cross sections given in (3.35) and (3.37). The mass-dependent contribution of the emitted particle, derived in App. F, is included and compared with the numerical integration of the full, unexpanded amplitude. To perform the comparison, we consider a system of distinguishable particles with equal coupling-to-mass ratios, ensuring that quadrupole emission provides the dominant contribution to the energy-differential cross section. Identical parameter values are used in both cases prior to any normalization, with fixed values of $v_{i},m_{j},m_{\phi,V}$, and the coupling constant $g_{j},y_{j},A_{j}/m_{j}$ and $\lambda_{j}$. The numerical results are in agreement with the analytical expressions, thereby validating the latter.
@@
Turning now to the case of a short-range interaction mediated by $\phi^{\prime}$ or $V^{\prime\mu}$, the quadrupole emission amplitudes remain model-dependent. In the limits $(r\ll 1,\,m_{\phi^{\prime},V^{\prime}}<\sqrt{s})$, the squared, averaged over the initials spins, summed over the final spins and averaged over the direction of the emitted particles amplitudes are obtained from (3.5.1) with the replacement $|\vec{v}_{i}-\vec{v}_{f}|^{4}\rightarrow m_{\phi^{\prime},V^{\prime}}^{4}/\mu^{4}$ and using the following quadrupole emission factors and coefficients for fermionic DM candidates instead,
@@
These emission amplitudes lead to the energy-differential cross sections
@@
In the case of scalar emission from scalar DM candidates, when the emission is vertex-dominated (see App. C), the squared emission amplitude is obtained from (3.36) via the substitutions $|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}\rightarrow m_{\phi^{\prime},V^{\prime}}^{4}/\mu^{4}$, $\lambda_{j}\to\lambda^{\prime}_{j}$, and $A_{j}\to A^{\prime}_{j}$. Similarly to the long-range mediation case, this squared emission amplitude does not originate from quadrupole emission, despite exhibiting the same $r$-dependence. The associated energy-differential cross section is
@@
We note that, in the contact limit, this energy-differential cross section depends only on the primed coupling of the DM to the short-range mediator and on an interaction involving both mediators simultaneously. In contrast, the quadrupole energy-differential cross sections (3.41) receive contributions from the primed couplings of the DM to the short-range mediator as well as from the unprimed couplings to the long-range mediator. For massive emission, the agreement between the $m_{\phi,V}$-dependent term of the $\phi$-emission amplitude and the longitudinal component of $V$-emission prior squaring again holds; see App. F for a comprehensive treatment. In the case of the energy-differential cross section (3.42), the $m_{\phi}$-dependent correction arises exclusively from the phase space.
In the left panel of Fig. 4, we show the energy-differential cross sections given in (3.41) and (3.42), including the mass-dependent contribution of the emitted particle, as derived in App. F. The latter are compared to a numerical integration of the full, unexpanded amplitude. To enable a direct comparison, we consider a system of distinguishable particles with equal coupling-to-mass ratios, for which quadrupole emission constitutes the leading contribution to the energy-differential cross section. Identical values of $v_{i}$, $m_{j}$, $m_{\phi,V}$, $m_{\phi^{\prime},V^{\prime}}$, and the coupling constants [^9] We use the following shorthand notation in the case of emission process involving scalar DM candidate: $C_{g^{\prime}_{j},\lambda_{S}}=g_{1}^{\prime}g^{\prime}_{2}+\lambda_{S}m^{2}_{V^{\prime}}/(4m_{1}m_{2})$, and $C_{A_{j}^{\prime},\lambda_{S}}=(A_{1}^{\prime}/m_{1})(A^{\prime}_{2}/m_{2})-\lambda_{S}m^{2}_{\phi^{\prime}}/(m_{1}m_{2})$. $g^{(\prime)}_{j}$, $y^{(\prime)}_{j}$, $A^{(\prime)}_{j}/m_{j}$, and $\lambda_{S}$ are used in both cases, prior to any normalization. We find excellent agreement between the analytical and numerical results.
@@ 3.5.2 Scattering of identical particles
For the emission of a vector boson $V^{\mu}$, in contrast to the scattering of distinguishable particles, the emission amplitudes of the dSIDM scenarios $\chi\chi\to\chi\chi V$ and $SS\to SSV$ are not identical. This difference arises from the statistical properties of the DM candidates, which determine the interference sign between the contributions from the $t$- and $u$-channel diagrams. For vector emission from Dirac fermions, $\chi\chi\to\chi\chi V$, the energy-differential cross section reads
@@
We now turn to the emission of $\phi$ particles from fermions and scalars. For scalar emission from fermions, both Dirac fermions, $\chi\chi\to\chi\chi\phi$ and Majorana fermions, $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$, share the same energy-differential cross section
@@
Here, we are in disagreement with the results of [83] where scalar emission from electron-electron and nucleon-nucleon bremsstrahlung is derived. We ascertain our analytical result by comparing it against an exact numerical evaluation of the full 2-to-3 cross section.
The energy-differential cross sections for $\phi$ emission from complex scalars, $SS\to SS\phi$, and real scalars $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$ are identical. In the limit $({A}/{m_{S}})^{2}\gg\lambda$, the external leg emission diagrams dominate, and the associated energy-differential cross section reads
@@
As a result, for fermionic DM candidates, the energy-differential cross section for massless emission is readily obtained by multiplying the right-hand side of (3.41) by the factor $S_{f}=1/2$, accounting for the symmetry of identical particles in the final state, and by performing the following substitutions: $g^{(\prime)}_{j}\to g^{(\prime)}$, $y^{(\prime)}_{j}\to y^{(\prime)}$, and $m_{j}\to m_{\chi}$. In the case of massive emission, these changes must be applied in (F.19) for the process $\chi\chi\to\chi\chi V$, and in (F.22) for the processes $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$ and $\chi\chi\to\chi\chi\phi$.
For scalar DM candidates, the inclusion of the $u$-channel doubles the contribution of the external leg and vertex diagrams to the emission amplitude compared to the contact emission diagrams. As a result, the dissipative process $SS\to SSV$ has the following energy-differential cross section
@@
The scenarios involving the emission of a scalar boson from real or complex scalar DM candidates, i.e., $SS\to SS\phi$, and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$, share identical energy-differential cross sections across the two limiting cases previously defined in Sec. 3.2. For external leg-dominated emission (i.e., $\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}\gg\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$; see App. C for details), we find
@@
For vertex-dominated emission (see App. C) in the scenarios $SS\to SS\phi$, and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$, the energy-differential cross section is
@@
Figure 4: Quadrupole energy-differential cross section for the different dSIDM scenarios mediated by a short-range force. The massless emission limits are shown as solid colored lines, while the dotted colored lines represent massive emission with $m_{\phi,V}/\omega_{\text{max}}=0.1$. The black stars indicate values of the energy-differential cross section obtained from a fully numerical integration of unexpanded amplitudes. The relevant coupling combinations $\{a,a^{\prime},a_{j}^{\prime}\}$ vary by scenario: i) for $\chi\chi\to\chi\chi V$, $\{g,g^{\prime},g_{j}^{\prime}\}$; ii) for $SS\to SSV$, $\{g,C^{1/2}_{g^{\prime},\lambda_{S}},C^{1/2}_{g_{j}^{\prime},\lambda_{S}}\}$; iii) for scalar emission from $\chi$($\tilde{\chi}$), $\{y,y^{\prime},y_{j}^{\prime}\}$; iv) for scalar emission with SS($\tilde{S}$) iv.a) in the external-leg regime (see App. C), $\{A/m_{S},C^{1/2}_{A^{\prime},\lambda_{S}},C^{1/2}_{A^{\prime}_{j},\lambda_{S}}\}$, iv.b) in the vertex-dominated regime (see App. C), $(a/m_{S,\chi})^{2}a^{\prime 2}_{1}a^{\prime 2}_{2}\to((A^{\prime}_{2}/m_{2})\lambda^{\prime}_{1}/m_{1}+(A^{\prime}_{1}/m_{1})\lambda^{\prime}_{2}/m_{2})^{2}$. Left panel: Distinguishable particle scattering (massless emission in Sec. 3.5.1 and massive emission in App. F) with equal coupling-to-mass ratio $a/m$. Numerical integration parameters: $m_{1}=0.1$ GeV, $m_{2}=1$ GeV, $a_{1}=0.1\,a_{2}$, $v_{i}\approx 4\times 10^{-3}$, and $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV. Right panel: Identical particle scattering (Sec. 3.5.2). Numerical integration parameters: $m_{S,\chi}=1$ GeV, $v_{i}\approx 6\times 10^{-3}$, and $m_{\phi^{\prime},V^{\prime}}=0.1$ GeV. Additionally, we assume $\lambda_{S}\sim g^{\prime}_{j},A^{\prime}_{j}/m_{j}$.